### v1

# Benchmarking Unified Face Attack Detection via Hierarchical Prompt Tuning

Abstract
Presentation Attack Detection (PAD) and Face Forgery Detection (FFD) are proposed to protect face data from physical media-based Presentation Attacks (PAs) and digital editing-based DeepFakes (DFs), respectively. However, isolated training of these two models significantly increases vulnerability towards unknown attacks, burdening deployment environments. The lack of a Unified Face Attack Detection (UAD) model to simultaneously handle attacks in these two categories is mainly attributed to two factors: (1) A benchmark that is sufficient for models to explore is lacking. Existing UAD datasets only contain limited attack types and samples, leading to the model’s confined ability to address abundant advanced threats. In light of these, through an explainable hierarchical way, we propose the most extensive and sophisticated collection of forgery techniques available to date, namely UniAttackDataPlus (UniAttackData+). Our UniAttackData+ encompasses $2,875$ identities and their $54$ kinds of corresponding falsified samples, in a total of $697,347$ videos. (2) The absence of a trustworthy classification criterion. Current methods endeavor to explore an arbitrary criterion within the same semantic space, which fails to exist when encountering diverse attacks. Thus, we present a novel Visual-Language Model-based Hierarchical Prompt Tuning Framework (HiPTune) that adaptively explores multiple classification criteria from different semantic spaces. Specifically, we construct a Visual Prompt Tree (VP-Tree) to explore various classification rules hierarchically. Then, by adaptively pruning the prompts, the model can select the most suitable prompts guiding the encoder to extract discriminative features at different levels in a coarse-to-fine manner. Finally, to help the model understand the classification criteria in visual space, we propose a Dynamically Prompt Integration (DPI) module to project the visual prompts to the text encoder to help obtain a more accurate semantics. Extensive experiments on 12 datasets have verified the potential to inspire further innovations in the UAD field.
Index Terms:
Unified Face Attack Detection, Face Anti-Spoofing, Face Forgery Detection, Prompt Tuning.
I Introduction
Face data are widely used in fields such as facial payment, mobile unlocking, and access authentication, due to the discriminative and easily obtainable facial biometric features. However, attackers can forge or steal identities in two ways to engage in illicit activities. As shown in the upper left of Fig. 2 (a), attackers can deceive face recognition systems by using physical media to impersonate others before information collection—a method known as physical attacks or presentation attacks. Additionally, attackers may employ digital forgery (commonly referred to as deepfake) to edit or manipulate collected images for fraudulent purposes. To address the aforementioned threats, Face Anti-Spoofing (FAS), also known as Presentation Attack Detection (PAD), and Face Forgery Detection (FFD) have been proposed to ensure the security of facial data. Essentially, both methods are formulated as binary classification tasks, where the model determines the authenticity of an image by outputting a probability indicating whether it is real or fake.
Researchers have proposed a wide variety of attack methods and have constructed multiple databases to advance the development of defense algorithms. Common physical presentation attacks are generally categorized into two types: 2D attacks (such as printed paper, fabric, screen replays) and 3D attacks (including 3D masks and silicone head models). Most models achieve authenticity classification by analyzing image texture details, focusing on inconsistent edges, screen moiré patterns, and other indicators. However, their generalization performance has been challenged. Consequently, domain generalization protocols such as ICMO and SCW have been proposed to evaluate model generalization. For instance, [1] utilized X, [2] leveraged X, and [3] employed X, while recent studies [4], [5], [6] have achieved notable results using vision-language models by employing X for X. Digital attack methods can be broadly classified into three categories: editing, adversarial attacks, and generative attacks. Digital editing attacks aim to modify image attributes (e.g., changing facial expressions, age) or perform face swapping (replacing the current face with another), video-driven manipulation (using a video to drive facial expressions), and more. Adversarial attacks can be divided into pixel-level and semantic-level attacks, which either modify or erase relevant pixels or alter high-dimensional semantics to manipulate the original facial data. Digital generative attacks, the most recent and innovative type, involve generating entirely new facial data using generative models. To combat digital attacks, researchers typically focus on detecting inconsistencies in pixel-level and semantic-level information to achieve robust classification performance.
Figure 1: The dashed lines represent the classification criteria that separate real and fake faces in the feature space. (a) Existing models are trained to draw a single ”line” in the feature space, leading to unpromising classification results. (b) Our HiPTune can adaptively integrate the classification criteria from different semantic spaces to achieve robust detection towards various attacks. Different colors represent distinct classification criteria.
Figure 2: (a) The Physical and Digital Attack construction approaches. Physical Presentation Attacks aim to steal identity information before the capture devices, and Digital Attacks attempt to edit the existing image to falsify the face data. (b) Our UniAttackData+ is constructed in an explainable hierarchical way. The coarse-to-fine categories provide researchers with prior knowledge which is helpful for the model designs. (b) A comparison of HiFAB with existing FAS, FFD, and UAD datasets shows that HiFAB surpasses all others in terms of attack types and video quantity. The radius of the circle represents the scale of the dataset, which is the number of videos contained in the dataset, corresponding to the horizontal axis of the figure. (Note: Grandfake is image-based) The vertical axis represents the number of attack types included in the dataset. (d) Sample demonstrations of the UniAttackData+. The samples are selected from the second level of the structure. The number following each category name indicates the number of attack types included.
Existing PAD and FFD models are typically trained and deployed separately, which inevitably reduces their robustness against unknown attacks and increases deployment complexity. When confronted with unknown attack samples or hybrid attack types, models trained independently exhibit poor robustness against such attacks. A unified model that can leverage the advantages of different attack types while maintaining a lightweight parameter size is urgently needed. We attribute the lack of such a unified model to two primary reasons:
(1) The lack of a comprehensive dataset that provides the model with a broader learning space. Although there are currently three unified face attack detection datasets (JFSFDB, Grandfake, UniAttack), they simply mix existing datasets that are tailored to specific attack types. This approach introduces substantial noise, such as variations in style, background, and identity, making it difficult to accurately evaluate the model’s superiority and generalization capability. Furthermore, as shown in Fig. 2 (b), their sizes are relatively small, containing limited attack types, and many of them are not publicly accessible, making them difficult to utilize effectively. To mitigate these issues, we previously proposed UniAttackData, a dataset where a total of 14 attacks (2 physical and 12 digital) are applied to the same face, creating an identity-consistent face attack dataset. However, its size and the diversity of attack types remain insufficient, resulting in models that tend to overfit the training set and perform poorly when encountering previously unseen attack types.
(2) The absence of a reliable classification criterion that can accurately categorize each sample. Existing methods typically aim to identify a single classification criterion within a single semantic space, as illustrated in Fig. 1 (a), where the model is trained to discover a single ’line’ that separates real and fake faces. However, in the context of UAD, this single classification criterion cannot robustly classify all sample types. From the perspective of data distribution, two major challenges arise: First, the diversity of attack types is constantly expanding, and these attacks increasingly resemble real samples across various dimensions. A single classification criterion is inadequate to address the rapidly evolving and sophisticated forgery techniques. Second, the significant feature differences between physical and digital attacks result in large intra-class variance among forged samples. The feature extractor struggles to effectively cluster them within a common space, making it difficult for a single classification criterion to determine the class of each sample accurately.
To address the dataset limitations mentioned above, we adopted the identity-consistent approach of UniAttackData and developed UniAttackDataPlus (UniAttackData+), which is currently the largest publicly available unified face attack detection dataset. It features the most extensive variety of attack types, minimal interference, and highly structured characteristics, making it a robust and comprehensive resource for training and evaluating face attack detection models. We collected data from $2,875$ participants representing three distinct ethnic groups (African, East Asian, and Central Asian), capturing $18,250$ authentic videos under various lighting conditions, backgrounds, and acquisition devices. For each participant, we applied $54$ different attack methods, including $14$ physical attacks and $40$ digital attacks, resulting in a total of $679,097$ forged videos, which is $32$ times more than the existing dataset. From each video, we randomly extracted at least $30$ frames to construct the dataset; samples can be found in Fig. 2 (c). The diverse acquisition backgrounds of authentic faces provide a more varied dataset, encouraging models to focus on discriminative features rather than irrelevant factors. Moreover, the extensive volume of data and the wide variety of attack types offer a more comprehensive and expansive learning space, motivating researchers to design more sophisticated and advanced models. Most importantly, UniAttackData+ applies a full range of attack methods to each unique identity, covering almost all known attack types. This not only prevents models from achieving high performance through ’identity memorization’ but also enhances the dataset’s applicability in real-world deployment scenarios. Additionally, our dataset is structured in a hierarchical manner, with all attack types categorized from coarse to fine, as depicted on the right side of Fig. 2 (a). This hierarchical structure provides researchers with a more interpretable framework for designing corresponding algorithms. For instance, our baseline method, HiPTune, leverages a similar hierarchical approach to develop a prompt-tuning method. Furthermore, we have designed four protocols to evaluate model performance under various scenarios, including unknown identities and unseen attack methods. We also introduced a cross-dataset testing benchmark to assess model generalization. Compared to previous datasets that simply merged existing datasets, UniAttackData+ offers a robust benchmark for the face data security community in a more interpretable hierarchical format, significantly promoting the development of more advanced algorithmic models.
To address the issue of insufficient robustness caused by a single classification criterion in unified attack detection, we propose a hierarchical prompt strategy named Hierarchical Prompt Tuning (HiPTune). HiPTune is designed to explore multiple classification criteria across different semantic spaces and adaptively integrate the most appropriate criterion for each sample. Specifically, following the hierarchical structure of our dataset (as shown on the right side of Fig. 2 (a)), we construct a Visual Prompt Tree (VP-Tree) as a carrier of multiple classification criteria. These learnable parameters, referred to as visual prompts, are concatenated with image tokens and fed into a fixed image encoder for training. These prompts serve as guiding symbols that direct the encoder to extract critical discriminative features, which we define as our ’classification criteria.’ The visual prompts are organized in a tree structure, progressing from coarse to fine, where each level represents a different granularity of classification criteria, and each leaf node corresponds to a specific criterion that is robust enough to distinguish certain fake samples from real faces. By inserting these visual prompts into different depths of the Transformer blocks in a top-down manner, we ensure that the model first learns to capture macro-level differences, achieving comprehensive detection of forgeries. As the feature extraction process deepens, finer-grained prompts focus on local details, enabling precise forgery discrimination. Shown as Fig. 1 (b), our HiPTune aims to utilize coarse-to-fine criteria to achieve a more adequate classification. This design endows the prompts with both global discriminative power and local specialization: top-level prompts enhance recall for unknown forgeries, while lower-level prompts improve fine-grained recognition of known categories. Moreover, to enrich the semantic information of these prompts, we introduce a Dynamic Prompt Integration (DPI) module, which employs a self-attention structure to adaptively integrate prompts from different levels and map them into the text embedding dimension. This cross-modal integration strategy broadens the semantic coverage of classification criteria, ensuring classification accuracy from multiple dimensions. Subsequently, we calculate the self-attention scores of the input samples, guiding the model to adaptively select the appropriate classification criteria at different levels for each sample. This adaptive strategy preserves model autonomy, avoiding excessive dependence on manual knowledge while uncovering potential connections between different attacks. It enables more robust detection tailored to each input sample.
To summarize, our contributions are as follows:
  • We create UniAttackData+, the largest publicly available unified face attack detection dataset, encompassing $54$ attacks per identity in an explainable hierarchical manner.
  • We proposed the Hierarchical Prompt Tuning (HiPTune) framework, which adaptively integrates classification criteria from different semantic spaces to establish a more robust criterion for each input sample.
  • We propose four intra-dataset protocols and nine cross-dataset benchmarks for the Unified Face Attack Detection community that promote the development of the detection model.
II Related Works
II-A Unified Face Attack Detection
II-A1 Datasets
Recently, several datasets, including Grandfake [1], JFSFDB [2], and UniAttack [3], have been introduced to address the challenges of UAD. These datasets combine existing PAD and FFD datasets to encompass a variety of attack types within a single dataset. Additionally, UniAttackData [4] was proposed as a UAD dataset with identity (ID) consistency, where each ID includes a complete range of attack types, thereby reducing ID and domain noise. The current UAD dataset not only lacks emerging generative forgery face samples but also outdated forgery algorithms corresponding to digitally edited faces, making it impossible to train a high-quality joint attack detection model. Our UniAttackData+ addresses this gap by incorporating a broader range of modern attacks and structuring them in an explainable manner, rather than merely aggregating datasets.
II-A2 Approaches
For the UAD task, Deb et al. [1] introduced a multi-task learning approach that treats each attack type as a separate detection task based on feature similarity. Yu et al. [2] focused on improving generalization by utilizing a multi-modal approach for joint task training. Fang et al. [4] leveraged the Vision-Language Model to capture both unified and specific knowledge, while Cao et al. [3] aimed to enhance face anti-spoofing by promoting the learning of compact, fundamental features of genuine faces. However, these methods rely on a single, generalized classification criterion, which often proves inadequate when faced with advanced attacks. In contrast, our HiPTune framework adaptively integrates the most suitable prompts from various semantic spaces, enabling more effective classification.
II-B Prompt Learning
PREFER [5] enhances prompt ensemble learning with a feedback mechanism for iterative refinement and prompt bagging, improving stability and performance across tasks. VPT [6] extends prompt learning from NLP to visual tasks. S-Prompt [7] clusters domain samples using K-NN, while C-Prompts [8] selects prompts from different domains for robust Domain Incremental Learning. CoOp [9] demonstrated that replacing manual prompts in CLIP with learnable contexts enhances performance via few-shot learning. CoCoOp [10] further adapts prompts to new domains with a Meta-Net, enhancing generalization.
However, existing methods overlook hierarchical prompt relationships, which capture both global and detailed knowledge. HiPro [11] organizes tasks hierarchically but lacks clarity between tasks. Our proposed HiPTune is able to adaptively select and integrate classification criteria from different semantic spaces by the VP-Tree, ensuring that the model not only learns to capture macro-level differences but also focuses on local details for the precise forgery discrimination.
Dataset Attack Type (each ID) # Datasets / Data # ID Physical Attacks Digital Attacks
Dataset Name No. # Categories No.
GrandFake Incomplete 6 sets: 789,412 (I) (Live: 341738, Fake: 447674) 96,817 SiW-M 128,112 (I) Adv (6) 116,641 (I)
DeepFake (6) 202,921 (I)
JFSFDB Incomplete 9 sets: 27,172 (V) (Live: 5,650, Fake: 21,522) 356 SiW 3,173 (V) DeepFake (4) 13,752 (V)
3DMAD 85 (V)
HKBU 588 (V)
MSU 210 (V)
3DMask 864 (V)
ROSE 2,850 (V)
UniAttackData Complete 1 set: 28,706 (V) (Live: 1,800, Fake: 26,906) 1,800 CASIA-SURF CeFA 5,400 (V) Adv (6) 10,706 (V)
DeepFake (6) 10,800 (V)
UniAttackData+ Complete 3 set: 697,347 (V) (Live: 18,250, Fake: 679,097) 2,875 CASIA-SURF 6,000 (V) Adv (16) 266,576 (V)
CeFA 9,000 (V) DeepFake (17) 242,859 (V)
HiFiMask 40,950 (V) Generation (7) 113,712 (V)
TABLE I: Comparison of existing UAD datasets. UniAttackData+ surpasses all other datasets in scale and diversity. ”V” and ”I” indicate video- and image-based counts, respectively. The numbers following each attack type in the ”# Catagories” row represent the number of algorithms included in that category.
Dataset # Attack Types # Videos
CASIA-SURF phy.6 21,000
CeFA phy.3 23,538
HiFiMask phy.5 54,600
SiW-v2 phy.14 1,700
MSU-MFSD phy.2 440
CASIA-FASD phy.3 600
ReplayAttack phy.2 1200
OULU-NPU phy.2 5940
FaceForensics++ digi.5 6,000
DeeperForensics-1.0 digi.1 60,000
DFDC digi.8 128,064
ForgeryNet digi.15 221,247
UniAttackData+ phy.10, digi.40 697,347
TABLE II: Comparison of existing PAD and FFD datasets. The scale of our UniAttackData+ is far more than the types in the existing PAD and FFD datasets. The ”#” represents ”the number of”.
III UniAttackDataPlus
III-A Advantages of UniAttackData+
Our UniAttackData+ dataset addresses five critical issues in current UAD datasets, i.e., insufficient sample size, limited attack coverage, and lack of interpretability, providing a comprehensive and valuable benchmark for the UAD community.
Scale. UniAttackData+ includes $2,875$ identities spanning three ethnicities, totaling $697,347$ video segments, with at least $25$ frames per video. Of these, $18,250$ are authentic face videos, 55,950 are physical attack videos, and $623,147$ are digital attack videos. Detailed data is presented in Tab. I and Tab. II. Our UniAttackData+ shows its superiority in separate PAD and FFD datasets, and also in the existing UAD datasets. The large dataset ensures a comprehensive and complete learning space, guaranteeing the effectiveness of detection models from a data perspective.
Attack Types. UniAttackData+ comprises $54$ attack types in total, including $14$ physical attacks and $40$ digital attacks, covering all known attack types in the field. This diversity ensures the model does not overfit to a specific attack type, thereby enhancing its generalization ability. All the attack methods included in our dataset were proposed after 2021, incorporating the latest digital generation attacks. UniAttackData+ ensures the data’s applicability in real-world scenarios, significantly expanding the range of attack types for model training and reducing vulnerability to emerging advanced attacks.
Interpretability. Existing UAD datasets merely merge different datasets without any correlation between the data, attack methods, and identities. This simplistic combination hinders the model from focusing on genuine forgery cues and greatly reduces model robustness in real-world applications. In contrast, UniAttackData+ preserves the identity consistency from UniAttackData while structuring all attack types hierarchically. Each attack method and sample is categorized from coarse to fine and labeled accordingly. This construction method makes full use of prior knowledge, offering researchers an actionable perspective to address UAD challenges. The hierarchical label collection and setup provide a more comprehensive experimental and analytical space, enabling deeper and more interpretable analyses of detection models and algorithms.
III-B Collection of UniAttackData+
Live Face source. The live faces in the UniAttackData+, enhancing diversity in terms of face identity, angle, expression, and ethnicity, are sourced from three public datasets: CASIA-SURF [12], which includes $1,000$ subjects; CASIA-SURF CeFA [13] (CeFA), comprising 1,800 subjects from three ethnicities - African, East Asian, and Central Asian, with $600$ subjects each ethnicity; and CASIA-SURF HiFiMask [14] (HiFiMask), which contains $75$ subjects. These varied sources significantly enhance the dataset’s diversity across multiple dimensions. We merge three existing datasets for the live face data to obtain $2,875$ distinct identities. For CASIA-SURF, the Intel RealSense SR300 camera captures collectors turning left or right, moving up or down, and approaching or moving away from the camera. The same collection method is used for the CASIA-SURF CeFA dataset. HiFiMask contains multiple lighting conditions, such as white, green, and periodic three-colors. Additionally, supplementary light sources from six directions are used to simulate real-world application scenarios.
For dataset preprocessing, we apply MTCNN [15] for face detection and cropping on the real face data from the original datasets, with the scale parameter set to 1.3. We then manually assign class labels, attack type labels, and hierarchical labels to facilitate research.
III-C Spoofing Approach
To ensure diversity in physical attack methods within the UniAttackData+ dataset, we incorporate $14$ physical face spoofing techniques, including $11$ 2D presentation attacks and $3$ 3D mask attacks. The 2D attacks consist of $6$ paper-cutout variations and $2$ print attack types, as well as $3$ video-replay attacks to enrich the dataset diversity further. For 3D mask attacks, UniAttackData+ includes high-fidelity masks made from three different materials: transparent, plaster, and resin. These masks are recorded under varied lighting conditions and with multiple imaging devices for added variation.
2D Attacks. We include two kinds of print attacks recorded in the outdoor and indoor environments. Moreover, three kinds of screen replay attacks from different devices are utilized. For the cutouts, the printed flat or curved face images will be cut into eyes, nose, mouth areas, or their combinations, generating 6 different attack ways. Specifically, the six ways are: (1) One person holds his/her flat face photo where eye regions are cut; (2) One person holds his/her curved face photo where eye regions are cut; (3) One person holds his/her flat face photo where eye and nose regions are cut; (4) One person holds his/her curved face photo where eye and nose regions are cut; (5) One person holds his/her flat face photo where eye, nose, and mouth regions are cut; (6) One person his/her curved face photo where eye, nose, and mouth regions are cut.
3D Attacks. For the 3D attacks, we construct three kinds of masks with different materials: transparent, plaster, and resin. All masks are worn on a real person’s face and uniformly dressed to avoid the algorithm looking for clues outside the head area.
Category Type Methods # Videos Total
2D Print Cloth-Indoor 1,800 3,600 55,950
Cloth-Outdoor 1,800
Replay pic-phone 1,800 5,400
screen-phone 1,800
screen-rs 1,800
Cutouts Cutout1 1,000 6,000
Cutout2 1,000
Cutout3 1,000
Cutout4 1,000
Cutout5 1,000
Cutout6 1,000
3D Transparent 13,650 40,950
Plaster 13,650
Resin 13,650
TABLE III: The detailed statistics of physical attacks in UniAttackData+. The ”#” represents ”the number of”.
III-D Forgery Approach
III-D1 Digital Manipulation Attacks
Digital Manipulation Attacks in UniAttackData+ encompasses $17$ specific methods grouped into three types of attacks: (1) Face Swapping [16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26], where identity features of a source face, $I_{\text{source}}$, are transferred into a target face, $I_{\text{target}}$, producing a composite image, $I_{\text{composite}}=f(I_{\text{source}},I_{\text{target}})$. (2) Facial Attribute Editing [27, 28], which alters specific attributes of a face, such as age, expression, or gender, expressed as $I^{\prime}=g(I,A_{\text{target}})$, where $A_{\text{target}}$ denotes the desired attribute. (3) Video-Driven Facial Animation [29, 30], which generates facial movements in static images using motion parameters $M$ from a driving source, represented by $I_{\text{animated}}=h(I,M)$.
Category Type Methods # Videos Total
Digital Manipulation Attribute-edit SAFA [27] 15,998 32,659 242,859
StarGAN [28] 16,661
Face-Swap FaceDancer [16] 16,168 176,878
GHOST [17] 13,135
SimSwap [18] 10,563
InsightFace [19] 13,264
BlendFace [20] 9,645
DeepLiveCam [31] 12,059
REFace [21] 12,756
FOMM [22] 16,661
InfoSwap [23] 16,534
MegaFS [24] 14,334
HifiFace [25] 16,661
DiffSwap [26] 13,761
Roop [32] 11,337
Video-Driven DaGAN [29] 16,661 33,322
OneShotTH [30] 16,661
TABLE IV: The detailed statistics of digital manipulation attacks.
Face-Swap. FaceDancer [16] utilizes Adaptive Feature Fusion Attention (AFFA) and Interpreted Feature Similarity Regularization (IFSR) to achieve seamless face-swapping with high-fidelity identity transfer and pose preservation. GHOST [17] introduces architectural enhancements like an eye-based loss function, face mask smoothing, and stabilization techniques to achieve high-fidelity image-to-image and image-to-video face swapping with improved identity preservation and gaze alignment. SimSwap [18] employs an ID Injection Module (IIM) for feature-level identity transfer and a Weak Feature Matching Loss to preserve target facial attributes, enabling high-fidelity and generalized face swapping. BlendFace [20] introduces a novel identity encoder trained on blended images to mitigate attribute biases, enabling better identity-attribute disentanglement for improved face-swapping while maintaining competitive performance. Deep-Live-Cam [31] provides tools for animating custom characters and modeling tasks, with built-in safeguards to prevent processing of inappropriate content, emphasizing ethical and lawful use, user responsibility, and adherence to legal requirements, including potential watermarking or project suspension if necessary. REFace [21] leverages diffusion models to frame face-swapping as a self-supervised inpainting task, incorporating DDIM sampling, CLIP feature disentanglement, and mask shuffling for enhanced identity transfer, realism, and versatility, including head and accessory swapping. FOMM [22] introduces unsupervised motion representations for animating articulated objects by identifying and tracking distinct parts, using region-based descriptions to disentangle shape and pose, and modeling global motion with affine transformations for improved animation fidelity. InfoSwap [23] uses an information bottleneck and identity contrastive loss to enhance identity disentanglement, generating high-quality, identity-discriminative swapped faces. MegaFS [24] introduces a Hierarchical Representation Face Encoder (HieRFE) and a Face Transfer Module (FTM) for high-quality megapixel-level face swapping, leveraging StyleGAN2 for stable synthesis and low memory usage. HifiFace [25] employs 3D shape-aware identity with geometric supervision and a Semantic Facial Fusion module for adaptive blending, achieving superior identity preservation and photo-realistic face-swapping results. DiffSwap [26] reframes face swapping as a conditional inpainting task using diffusion models, introducing a midpoint estimation method for efficient training and leveraging 3D-aware landmarks for controllable, high-fidelity, and shape-preserving face swapping. Roop [32] takes a video and replaces the face in it with a face of your choice. You only need one image of the desired face.
Attribute-edit. SAFA [27] employs 3D morphable models (3DMM) for facial structure modeling and inpainting techniques for occlusion recovery, enabling realistic and structure-aware face animation even for significantly deviated poses. StarGAN [28] employs a unified model architecture that enables multi-domain image-to-image translation within a single network, achieving superior scalability, robustness, and image quality compared to existing models.
Video-Driven. DaGAN [29] leverages self-supervised depth learning to recover dense 3D facial geometry, using it for keypoint estimation and 3D-aware cross-modal attention to guide motion field generation, enabling realistic and accurate talking head video synthesis. OneShotTH [30] uses a novel keypoint representation to decompose identity and motion information unsupervisedly, enabling high-quality talking-head video synthesis with reduced bandwidth and interactive head rotation for enhanced video conferencing experiences.
III-D2 Digital Adversarial Attacks
UniAttackData+ includes $16$ attacks designed to mislead machine learning models through subtle, often imperceptible perturbations. These attacks are categorized into pixel-level attacks [33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44], which modify raw image pixels to create adversarial examples, and semantic-level attacks [45, 46, 47, 48], which alter high-level features to change model interpretations. Pixel-level attacks rely on precise, gradient-based adjustments, while semantic-level attacks manipulate intermediate representations or latent spaces, increasing their transferability and effectiveness.
Category Type Methods # Videos Total
Digital Adversarial Pixel_level ALMA [33] 16,661 183,271 266,576
AdvDrop [34] 16,661
AI_FGTM [35] 16,661
ILA-DA [36] 16,661
AutoAttack [37] 16,661
PGD [38] 16,661
Jitter [40] 16,661
NIFGSM [41] 16,661
PIFGSMpp [42] 16,661
Pixel [43] 16,661
VMIFGSM [44] 16,661
Semantic-Level Demiguise [45] 16,661 83,305
SSAH [46] 16,661
DeCoWA [39] 16,661
PNA_PatchOut [47] 16,661
TGR [48] 16,661
TABLE V: The detailed statistics of digital adversarial attacks.
Pixel-Level. ALMA [33] utilizes Augmented Lagrangian-based attack for efficient, minimally perturbed adversarial examples, combining generality and speed. AdvDrop [34] introduces a novel adversarial attack by dropping imperceptible details from images instead of adding disturbances, challenging DNN robustness and evading existing defense mechanisms. AI-FGTM [35] improves adversarial example generation with dynamic step size, smaller kernels, and a tanh-based gradient method, achieving higher transferability, indistinguishability, and attack success rates than state-of-the-art methods. ILA-DA [36] enhances adversarial example transferability by combining automated image transformations, reverse adversarial updates, and attack interpolation, significantly outperforming prior methods in both defended and undefended scenarios. AutoAttack [37] introduces an ensemble of parameter-free adversarial attacks, improving PGD with optimized step size and objective function to rigorously evaluate adversarial robustness, uncovering weaknesses in many published defenses. PGD [38] explores adversarial robustness through robust optimization, introducing methods for training and attacking neural networks with universal security guarantees against first-order adversaries, significantly enhancing resistance to adversarial attacks. Jitter [40] proposes a novel loss function that improves attack efficiency and success rates by better exploring the perturbation space and addressing model confidence issues. NI-FGSM [41] leverages Nesterov accelerated gradients, and SIM, optimizing perturbations across scaled copies, to enhance the transferability of adversarial examples, achieving higher success rates on defense models. PIM++ [42] enhances adversarial transferability by using a patch-wise iterative method with an amplified step size, a projection kernel for gradient redistribution, and temperature adjustments to balance transferability and white-box attack performance. Pixel [43] introduces a black-box attack that rearranges a small number of pixels to achieve high attack success rates across models and datasets, requiring few iterations and minimal perceptual changes. VMI-FGSM [44] enhances gradient-based adversarial attacks by stabilizing update directions with gradient variance, significantly improving transferability.
Semantic-Level. Demiguise Attack [45] crafts unrestricted, photorealistic perturbations based on perceptual similarity, improving fooling rates, transferability, and robustness against defenses while maintaining human imperceptibility and simulating real-world conditions. SSAH [46] attacks semantic similarity in feature representations for cross-dataset adversarial generalization and employs a low-frequency constraint to generate imperceptible perturbations, achieving high transferability and perceptual similarity. DeCoWA [39] introduces Deformation-Constrained Warping (DeCoW) to enhance adversarial transferability across model genera by augmenting inputs with controlled elastic deformations, effectively attacking both CNNs and Transformers across various tasks. TGR [48] reduces gradient variance in Vision Transformer (ViT) blocks token-wise, generating more transferable adversarial samples and improving attack performance. PNA-PatchOut [47] combines Pay No Attention (PNA), skipping attention gradients, and PatchOut, optimizing random patch subsets, to enhance adversarial transferability across Vision Transformers (ViTs) and CNNs.
III-D3 Digital Generation Attacks
UniAttackData+ comprise $7$ types leveraging diffusion models to produce highly realistic synthetic content. These attacks are divided into three categories: Identity-Consistent Generation, Style Transfer, and Prompt-Driven Generation. Identity-Consistent Generation [49, 50, 51] seeks to maintain a subject’s identity across various contexts, modeled as $I_{\text{gen}}=G(I_{\text{id}},z)$, where $I_{\text{id}}$ is the identity input and $z$ represents variations such as pose or expression. Style Transfer, as in PhotoMaker [52], adapts identity features to different styles, represented as $I_{\text{styled}}=G(I_{\text{id}},\mathbf{S}_{\text{style}})$, to create visually diverse outputs. Prompt-Driven Generation [53] utilizes textual or image prompts to guide synthesis, expressed as $I_{\text{prompted}}=G(P_{\text{text}},P_{\text{img}})$.
ID-Consistant. ConsistentID [49] enhances personalized portrait generation with a multimodal facial prompt generator for fine-grained control and an ID-preservation network with facial attention localization, achieving superior ID consistency and precision from a single reference image. PuLID [50] introduces a tuning-free ID customization method for text-to-image generation, combining a Lightning T2I branch with contrastive alignment and ID losses to ensure high ID fidelity and minimal disruption to original image elements. InstantID [51] introduces a plug-and-play module with IdentityNet, leveraging strong semantic and weak spatial conditions to achieve high-fidelity, single-image-based identity personalization, seamlessly integrating with pre-trained diffusion models like SD1.5 and SDXL.
Style-Transfer. PhotoMaker [52] introduces a stacked ID embedding to encode multiple ID images into a unified representation, enabling efficient personalized text-to-image generation with high ID fidelity, text controllability, and speed, supported by an ID-oriented data construction pipeline.
Prompt-Based. IP-Adapter [53] introduces a decoupled cross-attention mechanism to enable lightweight image prompting for text-to-image diffusion models, achieving multimodal generation with only 22M parameters, while maintaining compatibility with text prompts, structural controls, and custom models.
Category Type Methods # Videos Total
Digital Generation ID-Consistent ConsistentID [49] 15,436 46,321 113,712
PuLID [50] 16,210
InstantID [51] 14,675
Style Transfer PhotoMaker [52] 18,192 18,192
Prompt Based IP-Adapter_kolors [53] 16,449 49,199
IP-Adapter_sdplusface [53] 16,449
IP-Adapter_sdpxlface [53] 16,301
TABLE VI: The detailed statistics of digital generation attacks.
III-E Protocol
ID Partition Protocol $P1$. Simply merging existing PAD and FFD datasets can cause models to focus on ID noise rather than discriminative features. To address this, we leverage the ID consistency of UniAttackData+. In this protocol, we train the model on a selected subset of identities, using the remaining identities for validation and testing. Each ID contains all attack types.
Cross-attack methods Protocol $P2$. In real-world deployment scenarios, a myriad of attack methods may emerge. Therefore, models should not only fit to seen attack methods but must also be capable of handling unseen attack methods. This protocol is designed to evaluate the model’s generalization ability across different attack methods. We adopted all identities for training and randomly selected 50% of attack methods as the training set, 20% as the validation set, and 30% as the test set. For instance, in the case of digital attacks, we used the first seven types of digital editing face swap attacks as the training set, two types for validation, and four types for testing, in order to assess the model’s performance on unseen face swap attacks.
Cross-attack types Protocol $P3$. In many cases, models are trained only on simpler attack types but are tested against more advanced and complex attack methods. This protocol aims to assess the model’s generalization ability from simpler to more sophisticated attacks. We trained the model using all identities, with 70% of physical 2D attacks, digital editing, and adversarial attacks from each identity as the training set, and 30% as the validation set. Physical 3D attacks and digital generative attacks were used as the test set, to evaluate the model’s generalization ability on unseen, more advanced attack types.
IV Hierarchical Prompt Tuning (HiPTune)
Figure 3: (a) The VP-Tree establishes a hierarchical prompt tree for all attack types, organized into three levels of grouping categories. It assigns a learnable prompt for each attack type and integrates the hierarchical prompts into the final output. (b) The APP employs a hierarchical prompt pruning mechanism to select the most promising prompt for each sample. (c) The DPI aims to aggregate all candidate prompts from different levels of VP-Tree and projects them into the text encoder for subsequent similarity computation.
IV-A Preliminaries
Vision-Language Models (VLM) like CLIP [54] utilize both text encoder $\mathcal{T}(\cdot)$ (like Transformer [55]) and image encoder $\mathcal{I}(\cdot)$ (like ViT [56]), where the input is the (text, image) pair, to leverages broader supervision. The outputs of both encoders are then dot-multiplied to obtain the probability score $p$:
$p(y=i\mid\bm{x})=\frac{\exp\left(\cos\left(\bm{w}_{\bm{i}},\bm{f}\right)/\tau\right)}{\sum_{j=1}^{K}\exp\left(\cos\left(\bm{w}_{\bm{j}},\bm{f}\right)/\tau\right)},$ (1)
where $\left\{\bm{w}_{i}\right\}_{i=1}^{K}$ denotes the set of weight vectors generated by the text encoder, and $\bm{f}$ represents the features extracted by the image encoder. $K$ is the number of classes, and $\tau$ refers to the temperature parameter learned by CLIP, $\cos(\cdot,\cdot)$ means the cosine similarity. Moreover, CoOp [57] explores prompt engineering in the VLM, constructing learnable prompts $\bm{t}=[\mathrm{V}]_{1}[\mathrm{~{}V}]_{2}\ldots[\mathrm{~{}V}]_{M}[\mathrm{CLASS}]$ and feeding to the fixed Text Encoder. The $[\mathrm{V}]_{m}(m\in\{1,\ldots,M\})$ is the vectors with the same dimension as word embeddings, and $M$ is a hyperparameter specifying the number of context tokens. The prediction probability is further computed as follows:
$p(y=i\mid\bm{x})=\frac{\exp\left(\cos\left(\mathcal{T}\left(\bm{t}_{i}\right),\bm{f}\right)/\tau\right)}{\sum_{j=1}^{K}\exp\left(\cos\left(\mathcal{T}\left(\bm{t}_{j}\right),\bm{f}\right)/\tau\right)}.$ (2)
Visual Prompt Tuning (VPT) aims to fine-tune the pre-trained models with only a few learnable parameters. Given the input image tokens $\bm{x}_{i},i\in{1,...,l}$ with length $l$, the trainable vectors $\bm{P}=\left\{\bm{p}^{k}\in\mathbb{R}^{d}\mid k\in\mathbb{N},1\leq k\leq p\right\}$, where d is the dimension of the patch embeddings and $p$ is the number of prompts, are cat with the original tokens and fed into the fixed encoders $Input=cat[\bm{x}\cdot\bm{P}]$.
IV-B Overview of HiPTune
We proposed a HiPTune that adopts hierarchical prompts to find multiple classification criteria from different semantic spaces and further integrates them, aiming to form more robust classification criteria for each input sample. Shown in Fig. 3, first, we construct the Visual Prompt Tree to gain the various semantics spaces tailored for each attack detection task. Then, given the input image $\bm{x}$, we adaptively select the most suitable prompt from each level of the VP-Tree by our Adaptive Prompt Pruning module. In addition, we integrate the prompt candidates from each selected level of the APP to describe the complete features of the input sample and feed it to the text encoder for better alignment.
IV-C Visual Prompt Tree (VP-Tree)
Define the levels. We can obtain a hierarchical relationship between each attack type based on the spoof and forgery types. Apart from the subjective categorization from the dataset, the features of different attacks demonstrate a more compact grouping relative to other categories while maintaining distinct independence within each category. For a VP-Tree with $N$ degrees, the level ${Tree}^{i}(i\in{1,2,...,N})$ has $n$ prompts, and $n$ may vary according to the set-ups. In our case, we define $3$-levels for the VP-Tree shown as Fig. 3 (a). The first level distributes the most obvious data as live, physical, and digital attack subtrees. The physical subtree is first separated into 2D and 3D attacks, which represents the second level of the VP-Tree. Print, replay, and cutouts are three key types for the level 3 prompts of a 2D attack. As for the 3D attacks, the masks’ material is always considered to be the spoof clues, so we define transparent, plaster, and resin prompts within the 3D attack prompt. Concerning the digital subtree, we categorize the digital attacks into $3$ kinds according to their forgery techniques. The digital manipulation method includes attribute-editing, face-swap, and video-driven faces. Adversarial attacks are divided into pixel-level and semantic-level attacks for forgery areas. We also define $3$ sub-prompts for the digital generation attacks: ID-consistent generation, style transfer, and prompt-based editing forgeries.
Initialize the prompts. After we obtain the VP-Tree’s structure, we define the prompts for each level with random initialization in the following form:
$\bm{p}=[\mathrm{V}]_{1}[\mathrm{~{}V}]_{2}\ldots[\mathrm{~{}V}]_{l},$ (3)
where each $[\mathrm{V}]_{l}(l\in\{1,\ldots,L\})$ is the learnable vectors with the length of $L$ at the same dimension with the patch embeddings (i.e., 768 for CLIP image encoder).
Train and integrate the prompts. In order to force each prompt to learn the classification criteria specific to each attack type, we train the prompts in a supervised manner during the first stage. Given the input $\bm{x}\in\mathbb{R}^{C\times H\times W}$, where $C,H,W$ represent the channel, height, and width of the input images, we manually record the attack types and select the corresponding prompts at each level accordingly. For the live face images, since we train our model with (live, fake) pairs, we choose the same prompt as the partner. Then, we average the selected prompts to form the final prompt and link with the image tokens. The final input of the image encoder is formulated as Eq. 4:
${\mathcal{I}}_{input}=cat[(\bm{x}_{cls},\bm{x}_{1},...,\bm{x}_{l})\cdot\bm{p}],$ (4)
where $(\bm{x}_{cls},\bm{x}_{1},...,\bm{x}_{l})\in\mathbb{R}^{l+1,dim}$ is the image tokens after the patch embeddings. $l$ represents the length of image tokens, and $dim=768$ is the embedding dimension. The $\bm{p}$ is the final output of the integrated prompts as Eq. 5
$\bm{p}=mean(\bm{p}_{1}+\bm{p}_{2}+\bm{p}_{3}),$ (5)
where $\bm{p}_{1},\bm{p}_{2},\bm{p}_{3}$ are the selected prompts at level $1,2,3$ and $mean(\cdot)$ represents the average of the inputs.
After fully training the VP-Tree to represent the classification criteria for multiple semantic spaces, the Adaptive Prompt Pruning (APP) module will learn to select the most suitable prompts for each sample.
IV-D Adaptive Prompt Pruning (APP)
As shown Fig. 3 (b), aiming to explore classification criteria from different semantic spaces, the APP module adopts a hierarchical pruning mechanism to select the most suitable prompt for each sample. Given the input image $\bm{x}\in\mathbb{R}^{C\times H\times W}$, the patch embedding first projects the image into the visual embedding dimension and reshapes $\bm{x}$ into a series of image token with length $L$: $\bm{x}:\mathbb{R}^{C\times H\times W}\rightarrow\mathbb{R}^{L\times D}$. In our case, $C=3$ represents the channels for the image, and $H,W$ are both $244$ for the input image, the token’s length $L=196$, and the dimension of visual embedding is $D=768$. After adding the positional embeddings, a class token is linked with the image token, and then the image token $\bm{x}\in\mathbb{R}^{197\times 768}$ is fed to the APP.
Level-1 prompt pruning. Given the significant distribution variance between physical and digital attacks, it is simple to determine whether physical or digital criteria should be used to classify this sample. Thus, we utilize a simple linear layer to select the level 1 prompt:
$G_{1}(\bm{x})=\mathbf{W}_{1}\cdot\bm{x}+bias_{1},$ (6)
where $G_{1}$ is the gate function of the level-1 prompt pruning and $\mathbf{W}_{1},bias_{1}$ are learnable weights and bias. The level-1 router first projects each patch’s features into three alternatives using $G_{1}$ and then calculates the mean value over all tokens. After applying Softmax between three classes, we choose the $TOP_{K=1}$ candidate prompt. The whole procedure can be described as:
$i^{p}_{1}=TOP_{k=1}(softmax(mean(G_{1}(\bm{x}),197))),$ (7)
where $i^{p}_{1}$ denotes the index of the selected prompt, $mean(\cdot,\cdot)$ represents the mean calculus over specific dimension. Furthermore, the selected prompt is linked with the original image tokens and forwarded to the next level.
Level-2 prompt pruning. Regarding the level-2 prompt pruning, it is challenging to separate samples from attack types within each category if we only rely on fixed visual embeddings. In light of this, we apply Central Difference Convolution (CDC) [58] and the Frequency-Adaptive Dilated Convolution (FADC) [59] for level-2 prompt pruning in the physical and digital branches, respectively.
After the CDC or the FADC, the output feature map is sent to a linear layer, and then the tokens’ scores are averaged to select the candidate prompt further. The level-2 gate functions can be expressed as:
$G_{2}^{0}$ $=\mathbf{W}_{2}^{0}\cdot CDC_{2}^{0}(\bm{x})+bias_{2}^{0}$ (8)
$G_{2}^{1}$ $=\mathbf{W}_{2}^{1}\cdot FADC_{2}^{1}(\bm{x})+bias_{2}^{1},$
and the selected prompt can be denoted as:
$i^{p}_{2}=TOP_{k=1}(softmax(mean(G_{2}(\bm{x}),197))).$ (9)
The level-1 and level-2 prompts are combined averagely and concat with the image tokens and fed to the next router.
Level-3 prompt pruning. The level-3 prompt pruning procedure is the same as the level-2 pruning process. The gate functions can be described as follows:
$G_{3}^{i}=\mathbf{W}_{3}^{i}\cdot Conv_{3}^{i}(\bm{x})+bias_{3}^{i},$ (10)
where $i\in{1,2,3}$ for physical 2D/3D, digital manipulation, generation, and $i\in{1,2}$ for digital adversarial attacks.
Finally, the selected prompts from all three levels are averaged together and concated with the original image tokens. For the live images, the live prompt from the first level will be directly concated with the tokens and fed to the image encoder.
IV-E Dynamic Prompt Interaction (DPI)
CLIP’s default templates,  e.g., “This is a photo of [CLASS]”, fail to describe the detailed semantic information for the multiple attack types, which caused the performance drops encountering advanced attacks. Moreover, the unified prompts in the CoOp still can not represent the differences between various attacks and the live faces. However, as mentioned above, our VP-Tree constructs a hierarchical prompt tree for the different classification criteria, and our APP obtains the probabilities of the suitable candidate prompts. Thus, we proposed a Dynamic Prompt Interaction (DPI) to project the visual prompts to the text encoder to fully describe the abundant features corresponding to each category.
Specifically, first, we directly pass the live prompts from the first level as the live class text prompts. As for the fake prompts, we apply the probabilities that the app has predicted to the candidate prompts and combine them weightedly. For the first level, we apply the softmax function only to physical and digital prompt scores to obtain the binary probabilities of two categories. Then, we weighted sum the prompts to get a weighted level-1 fake prompt. In the second level, due to the router from the first level choosing a subtree for the sample and needing a possible fake prompt specific to the input sample, we repeat the weighted sum of candidate prompts until the router has not predicted the further probabilities. Lastly, we form prompts for both live and fake faces and send them to the text encoder for further similarity computation.
IV-F Objectives and Inference
The training objective for the first training stage is to help the prompts to learn the semantic information and classification criteria specific to the attack type. The second stage aims to train the APP to select the most suitable prompts tailored for each sample successfully. The loss for both stages is:
$\mathcal{L}=\mathcal{L}_{CE}+\mathcal{L}_{triple},$ (11)
where $\mathcal{L}_{CE}$ represents the cross-entropy loss and $\mathcal{L}_{triple}$ is the Asymmetric Triplet Loss [60]. The reason for applying Asymmetric Triplet Loss is to separate the different prompts to preserve their unique classification criteria while combat the relatively similar prompts.
V Experiments
V-A Experiment Settings
Datasets, Protocols, and Evaluation Metrics. We evaluated the effectiveness of the proposed UniAttackData+ dataset and the robustness of the HiPTune algorithm using three existing UAD datasets. For JFSFDB, we conducted both internal and external testing within its original FAS and FFD datasets, as well as within the joint training dataset. For UniAttackData, we tested using its original p1, p2.1, and p2.2 subsets. For our proposed UniAttackData+, we performed training and testing based on the four protocols we introduced. Additionally, we established a UAD-Benchmark to assess the model’s performance in cross-dataset testing. We evaluate model performance with four metrics: (1) ACER (average of FRR and FAR), (2) AUC (theoretical model performance), (3) EER (optimal threshold), and (4) ACC (overall accuracy).
Implementation Details. We use a $3$-level VP-Tree with a prompt length of $60$, a $\theta$-value of CDC set to $0.7$, and a fixed CLIP model with ViT/b-16 backbone, classifying ”live” and ”fake.” The initial learning rate is $10^{-5}$ with the Adam optimizer. Images are resized to $224\times 224$ with random cropping and flipping. VP-Tree training runs on an NVIDIA GeForce RTX 4090 for $100$ epochs (batch size: $32$), followed by full model training for $200$ epochs (batch size: $32$).
V-B Results on UniAttackData+
p1-ID p2-cross_attack_methods p3.1-cross_attacks(low2high) p3.2-cross_attacks(high2low)
Model ACER AUC ACC EER ACER AUC ACC EER ACER AUC ACC EER ACER AUC ACC EER
CLIP-V 9.36 92.46 88.05 9.36 28.77 71.99 78.49 28.77 49.87 43.12 58.08 49.87 51.17 47.35 86.76 51.17
CLIP 9.93 95.22 83.75 9.93 22.89 72.79 76.15 22.89 34.37 58.09 59.9 34.37 53.05 46.95 82.68 53.05
CoOp(unified) 14.76 90.89 83.92 14.76 43.14 59.64 47.83 43.14 72 20.89 27.18 72 45.06 53.7 59.2 45.06
CoOp(specific) 15.26 90.8 84.48 15.26 44.73 55.44 48.5 44.73 71.57 21.76 24.79 71.57 42.82 57.26 57.83 42.82
CDCN++ 12.77 94.53 90.89 11.72 15.23 91.78 84.72 15.22 37.55 67.9 58.67 38.46 42.09 62.16 70.55 39.66
CFPL-FAS 13.93 93.77 96.235 11.65 16.1 86.34 83.09 16.1 29.6 73.5 71.46 29.6 53.2 48 84.98 50.21
MoE-FFD 17.3 91.9 96.5 15.69 30.53 70.09 92.82 31.39 46.01 51.27 89.14 47.97 54.85 45.11 88.81 54.85
LLA-Net 10.57 95.64 95.97 7.61 18.34 82.97 89.86 21.31 36.58 57.02 68.25 42.2 54.04 45.94 87.71 54.04
FA-CLIP 8.26 93.96 94.53 8.45 35.81 63.89 88.51 47.37 44.44 48.34 56.95 41.05 51.22 48.78 90.38 99.64
HiPTune 7.31 97.87 90.02 7.61 14.26 92.5 90.26 14.71 25.68 77.62 69.34 26.02 40.13 65.89 92.66 40.18
TABLE VII: Results on UniAttackData+. CoOp(unified/specific) represents the CoOp without/with the class-specific context.
We employed a range of CLIP-based baseline models (including CLIP-V, CLIP, CoOp without class-specific context, and CoOp with class-specific context) to evaluate the performance of shared-space learning models on our dataset. In addition, we assessed algorithms specifically designed for PAD/FFD, such as CFPL-FAS, MoE-FFD, and LLA-Net, to determine their generalization capabilities. We also compared our approach with the recently released $FA^{3}-CLIP$, an algorithm specifically designed for unified attack detection. The detailed experimental results are presented in Table VII.
Authenticity-Only Testing Protocol $p1$:
This protocol primarily aims to evaluate the model’s true capability in distinguishing between real and fake samples. As shown in the experimental results table (Table VII), models that attempt to update the encoder generally achieve better performance. This is because, within UniAttackData+, a shared feature space still exists where different attack types can be clustered, and enhancing the image encoder’s feature extraction capability improves model performance. However, these models have a large number of parameters, making them unsuitable for deployment in real-world environments.
Conversely, CoOp, which relies on prompt matching, attempts to map fixed features to relevant regions. However, given the vast and unpredictable variety of attack types, a single prompt is insufficient to address the diversity and sophistication of these attacks. Models specifically designed for PAD and FFD focus solely on the manipulated regions of the respective images, neglecting the common forgery cues and underlying processes shared across UAD tasks. Notably, CDCN++, a purely convolutional model, also demonstrates decent performance, highlighting the importance of local region analysis in UAD tasks. This suggests that local feature extraction remains crucial for robust attack detection.
Our HiPTune adopts a visual prompt approach, which not only guides the encoder in feature extraction but also maintains model efficiency. Furthermore, our hierarchical prompt tree encapsulates multi-level, fine-grained classification criteria, achieving highly precise classification performance.
Real-World Robustness Testing Protocol $p2$:
This protocol is widely applicable in real-world scenarios, aiming to evaluate the model’s performance on seen attack types (such as physical 2D attacks) and unseen attack methods (such as face-swapping techniques like Roop and SimSwap). Similar to protocol p1, models that train the encoder generally achieve superior performance. This is because identifying a shared feature space helps the model generate a general mapping for previously unseen methods, effectively leveraging existing knowledge to handle new attack techniques.
In contrast, models like CoOp, which rely on prompt matching, are more prone to overfitting the training data. As a result, their detection performance significantly declines when encountering new data distributions. For instance, CDCN++, a model specifically designed for PAD, utilizes a central difference convolutional structure that emphasizes local detail information, achieving the best results in our benchmark experiments. This finding highlights the importance of focusing on local detail information to ensure model generalization in UAD tasks.
Our HiPTune model adopts a visual prompt approach, introducing prompts on the image encoder side to guide the frozen encoder in extracting relevant features while maintaining efficiency. Unlike conventional prompt-matching methods, our prompts serve only as guidance, avoiding excessive dependence on existing knowledge for decision-making. This design not only enhances the model’s generalization capability against unseen attack methods but also ensures robustness and interpretability.
Upward Compatibility Protocol $p3.1$:
This protocol aims to assess whether models trained on simple attack types can maintain high robustness in testing environments featuring complex and advanced attacks. We observe a sharp performance decline in both CoOp models, which we attribute to the absence of encoder guidance. Relying solely on text-side mapping relationships fails to achieve strong reasoning and generalization capabilities.
In contrast, our HiPTune model, which adaptively integrates different types of prompts, demonstrates superior performance. Despite being trained on limited samples, HiPTune can make judgments based on existing higher-level prompts while leveraging data from other attack types to assist in inference. This design ensures that the model retains strong generalization even in the presence of advanced attack types.
Downward Compatibility Protocol $p3.2$:
This protocol aims to evaluate whether models can leverage advanced knowledge to accurately detect simpler attack types. As shown in the table, most models exhibit poor performance, indicating that existing methods are insufficient to effectively utilize anti-forgery knowledge derived from advanced digital generative attacks. This also reveals the potential non-transferability between these two types of knowledge.
In contrast, our HiPTune model demonstrates superior performance. It effectively borrows knowledge from higher-level prompts and dynamically combines existing classification criteria, highlighting its advanced design. This adaptive approach allows HiPTune to generalize knowledge from complex attacks to simpler ones, ensuring robust performance across diverse scenarios.
V-C Results on UAD-Benchmark
JFSFDB-UniAttackData+ UniAttackData-UniAttackData+ JFSFDB+UniAttackData-UniAttacData+
ACER AUC ACC EER ACER AUC ACC EER ACER AUC ACC EER
CLIP-V 35.83 67.91 53.44 35.83 41.41 60.9 49.69 41.41 39.81 63.4 54.1 39.81
CLIP 39.32 63.95 55.65 39.32 40.58 62.75 50.38 40.58 39.36 64.08 51.92 39.36
CoOp(unified) 39.59 63.56 58.29 39.6 26.93 78.77 70.44 26.93 29.77 76.77 67.4 29.77
CoOp(specific) 39.39 63.89 58.71 39.39 28.62 77.5 69.9 28.62 30.3 75.9 68.49 30.3
HiPTune 32.88 69.34 60.51 32.88 20.16 88.53 86.82 20.23 25.93 83.05 73.82 25.94
UniAttackData-JFSFDB UniAttackData+-JFSFDB UniAttackData+UniAttacData+-JFSFDB
ACER AUC ACC EER ACER AUC ACC EER ACER AUC ACC EER
CLIP-V 36.16 64.97 70.81 36.16 51.65 48.75 47.88 51.65 45.12 56.26 50.52 45.12
CLIP 36.98 64.05 76.04 36.98 50.72 48.25 44.09 50.72 44.42 58.1 54.11 44.42
CoOp(unified) 40.02 64.12 58.53 40.02 43.55 59.51 54.28 43.55 38.21 65.84 59.44 38.21
CoOp(specific) 39.44 64.83 59.02 39.44 42.95 59.99 55.39 42.95 38.13 66.04 60.67 38.13
HiPTune 34.39 71.23 80.31 34.42 28.14 79.22 70.46 28.78 32.93 70.89 72.23 32.92
JFSFDB-UniAttackData UniAttackData+-UniAttackData JFSFDB+UniAttackData+-UniAttacData
ACER AUC ACC EER ACER AUC ACC EER ACER AUC ACC EER
CLIP-V 29.94 75.58 62.11 29.94 22.26 84.22 56.92 22.26 13.15 91.86 77.74 13.15
CLIP 29.27 80.12 67.63 29.27 22.38 84.11 75.76 22.38 16.59 90.14 82.5 16.59
CoOp(unified) 37.17 67.2 58.91 37.17 25.4 80.5 71.77 25.4 30.28 76.11 65.41 30.28
CoOp(specific) 41.4 61.95 57.39 41.4 26.37 79.32 70.68 26.37 27.34 79.19 67.3 27.34
HiPTune 21.36 85.59 74.39 31.56 18.62 90.75 88.64 18.63 8.55 94.67 90.26 8.56
TABLE VIII: Results on existing UAD datasets.
In the UAD-Benchmark, training and testing models across different datasets allow us to evaluate not only their generalization to various attack types but also their domain generalization capabilities. By examining the table, we can draw two key insights:
(1) As shown in the second column of the table, models trained on UniAttackData+ demonstrate strong anti-forgery capabilities that can be effectively transferred to other datasets. This highlights the robustness and generalizability of the knowledge learned from UniAttackData+.
(2) Baseline models show limited improvement, suggesting that we should design models specifically tailored to the characteristics of the dataset. Such models should be capable of fully leveraging the knowledge contained within the dataset, maximizing its utility for model training and evaluation.
V-D Results on JFSFDB
Face Spoofing Detection Face Forgery Detection
Intra-testing Cross-testing Intra-testing Cross-testing
Metrics Sep./ Joint Method SiW 3DMAD HKBU MSU 3DMask ROSE FF++ DFDC Celeb-DFv2 Average
MesoNet 95.73 100 68.87 85.33 44.54 68.05 65.54 54.72 65.1 71.99
Xception 99.88 100 94.47 90.62 63.88 90.15 98.21 64.86 84.27 87.37
MultiAtten 99.56 100 68.68 89.21 36.88 72.66 70.24 63.37 64.34 73.88
CDCN++ 100 99.56 94.51 82.65 47.66 76.76 97.53 56.65 78.34 81.51
DeepPixel 100 100 89.4 95.17 56.87 80.6 99.26 67.54 85.51 86.04
ResNet50 99.99 100 93.71 94.33 58.95 88.79 97.04 62.35 82.51 86.41
ViT 100 100 95.45 95.13 35.6 89.12 98.65 73.79 86.61 86.04
CLIP-V 99.99 100 96.85 97.29 76.62 93.6 93.05 82.09 75.4 90.54
CLIP 99.99 100 95.58 95.82 80.4 92.89 95.92 86.08 79.4 91.79
CoOp(Unified) 99.86 100 94.42 83.73 91.94 94.63 79.49 64.46 62.05 85.62
CoOp(Specific) 99.83 100 93.72 84.14 92.61 92.89 81.41 66.63 61.88 85.9
Sep. HipTune 99.97 100 98.62 98.62 95.87 98/62 100 89.72 89.63 93.12
MesoNet 92.55 97.94 78.42 84.01 45.87 70 56.66 50.57 55.03 72.56
Xception 99.8 100 97.62 91.56 75.49 79.38 96.69 63.69 78.87 87.01
MultiAtten 99.38 100 78.34 92.89 58.31 77.51 69 58.11 60 77.06
CDCN++ 99.99 98.89 96.8 78.93 61.54 76.39 95.79 56.42 76.92 82.41
DeepPixel 100 100 97.31 93.97 66.65 74.58 99.22 66.36 88.45 87.39
ResNet50 99.93 100 98.1 93.5 79.64 86.03 98.83 67.18 85.55 89.86
ViT 100 100 100 95.85 62.74 91.88 98.83 75.11 88.36 90.31
JFSFDB-Uni 100 100 99.96 97.99 67.1 95.57 99.12 75.9 89.76 91.71
CLIP-V 100 100 95.91 97.23 71.26 83.96 95.69 82.84 79.64 89.44
CLIP 100 100 97.48 96.39 86.13 81.07 94.83 84.06 78.99 90.83
CoOp(Unified) 99.71 100 95.82 81.58 89.27 93.62 78.55 68.56 64.88 85.78
CoOp(Specific) 99.78 100 95.33 80.68 92.82 93.34 76.42 67.58 65.06 85.67
AUC(%) Joint HipTune 100 100 100 99.62 95.43 97.82 100 90.88 93.35 97.46
TABLE IX: Results on JFSFDB AUC.
Face Spoofing Detection Face Forgery Detection
Intra-testing Cross-testing Intra-testing Cross-testing
Metrics Sep./ Joint Method SiW 3DMAD HKBU MSU 3DMask ROSE FF++ DFDC Celeb-DFv2 Average
MesoNet 11.67 0 35.12 21.9 55.82 36.81 38.93 46.16 38.78 31.69
Xception 1.14 0 11.31 18.57 42.3 18.13 5.54 38.92 24.07 17.78
MultiAtten 2.78 0 35.71 18.57 61.22 33.82 35.54 39.73 40.46 29.76
CDCN++ 0.14 3.33 11.9 22.8 52.76 32.01 7.5 44.75 28.19 22.6
DeepPixel 0 0 17.86 15.71 46.42 28.23 2.32 36.82 22.81 18.91
ResNet50 0.36 0 13.69 12.38 45.24 19.94 8.18 40.61 26.01 18.49
ViT 0.14 0 11.31 11.9 60.4 19.47 4.64 32.75 22.76 18.15
CLIP-V 0.49 0 7.74 8.58 29.56 13.17 13.66 25.43 29.89 14.28
CLIP 0.34 0 11.51 12.86 26.74 14.61 11.47 23.14 27.16 14.2
CoOp(Unified) 1.51 0 12.7 25.95 17.02 12.06 27.14 39.42 41.21 19.67
CoOp(Specific) 1.66 0 15.08 25.48 15.68 15.25 25.98 37.69 41.53 19.82
Sep. HipTune 0.35 0 4.58 6.56 11.57 12.82 1.89 20.86 28.77 9.71
Mesonet 14.8 6.67 28.57 22.86 54.76 36.18 44.29 49.81 46.94 33.88
Xception 1.71 0 9.52 17.14 22.21 28.82 9.29 40.75 29.16 17.62
MultiAtten 4.13 0 28.57 13.81 44.42 30.32 35.96 43.93 43.57 27.19
CDCN++ 0.28 6.67 10.71 31.9 43.83 33.19 8.39 44.58 29.68 23.25
DeepPixel 0.14 0 10.12 13.33 40.19 32.8 2.14 37.38 19.48 17.29
ResNet50 1.21 0 6.55 12.86 28.79 21.55 4.29 36.82 23.06 15.01
ViT 0 0 0 9.52 46.3 16.2 2.68 30.99 20.01 13.97
JFSFDB-Uni 0 0 0.6 6.67 43.6 11.64 2.32 30.6 18.8 12.69
CLIP-V 0 0 10.71 8.33 37.17 23.46 11.43 25.72 27.81 16.8
CLIP 0 0 8.13 10.24 23.58 26.86 12.14 23.76 27.34 15.4
CoOp(Unified) 0.02 0 11.11 28.57 17.77 14.39 28.57 36.69 39.84 19.66
CoOp(Specific) 1.54 0 11.11 29.76 13.82 14.48 31.43 37.22 40.31 19.96
EER(%) Joint HipTune 0 0 0 3.84 11.72 8.99 1.23 24.92 12.21 6.99
TABLE X: Results on JFSFDB-EER.
JFSFDB primarily demonstrated that training with a combination of physical and digital attack samples can significantly enhance model performance. This conclusion is also applicable to our HiPTune model. When trained and tested separately within the dataset, HiPTune achieved an EER of 9.71% and an AUC of 93.12%. However, after joint training, HiPTune gained access to a more diverse set of classification criteria, which it could adaptively select and integrate, resulting in a substantial performance improvement: a 4.34% increase in AUC and a 2.72% reduction in EER. Moreover, the model’s generalization capability was significantly enhanced.
V-E Results on UniAttackData
Protocol Methods ACER ACC AUC EER Protocol Methods ACER ACC AUC EER
ResNet-50 1.35 98.83 99.79 1.18 ResNet-50 34.6 53.69 87.89 19.48
ViT-B/16 5.92 92.29 97 9.14 ViT-B/16 33.69 52.43 83.77 25.94
Auxiliary 1.13 98.68 99.82 1.23 Auxiliary 42.98 37.71 76.27 32.66
CDCN 1.4 98.57 99.52 1.42 CDCN 34.33 53.1 77.46 29.17
FFD 2.01 97.97 99.57 2.01 FFD 44.2 40.43 80.97 26.18
CLIP-V 0.46 99.55 99.95 0.46 CLIP-V 11.38 85.93 95.73 11.38
CLIP 0.58 97.4 99.93 0.58 CLIP 15.27 79.83 92.46 15.27
CoOp(unified) 7.94 90.96 97.9 7.94 CoOp(unified) 33.29 63.89 71.58 33.29
CoOp(specific) 8.49 90.98 97.45 8.49 CoOp(specific) 35.67 60.9 68.41 35.67
UniAttackDetection 0.52 99.45 99.96 0.53 UniAttackDetection 22.42 67.35 91.97 15.72
1 HiPTune 0.33 99.98 99.99 0.36 2 HiPTune 10.38 88.62 97.82 10.39
TABLE XI: Results on UniAttackData.
In the UniAttackData dataset, where the data distribution is relatively simple, our HiPTune model achieved outstanding performance. Leveraging the guidance function of prompts, HiPTune also delivered excellent results in cross-attack protocols, demonstrating its robustness and adaptability across different attack types.
V-F Abalation Study
To verify the effectiveness of the theory proposed by HiPTune and to analyze the influence of various settings, we conduct several ablation studies on protocol $P_{2}$ of HiFAB. These studies examine the depth of the tree, length of prompts, different selection strategies, various gate functions, and the contribution of each component. Additional experiments and detailed analyses of each element are provided in the Supplementary Materials.
Depth of Tree. The depth of the VP-Tree plays a crucial role in helping models understand fine-grained classification criteria, which enhances their ability to handle unseen advanced attacks during the integration process. Models such as CoOp, which lack a hierarchical structure, exhibit poor performance as they rely on a single, rigid classification criterion. In contrast, increasing the depth of the VP-Tree allows the model to ”draw” the line between live and fake faces adaptively and accurately, demonstrating the importance of a hierarchical approach.
Length of Prompts. The length of prompts determines the scope of learnable parameters, impacting both accuracy and efficiency. To explore this balance, we conduct several experiments to assess the effect of different prompt lengths. Using more parameters can capture more detailed classification-related information, enhancing model performance.
Component ACER(%) AUC(%) EER(%) ACC(%)
level-0 (CoOp) 14.20 92.57 14.26 85.44
level-1 9.78 94,26 10.12 90.12
level-2 5.78 96.88 5.45 95.08
level-3 3.26 98.78 3.67 95.79
TABLE XII: The influence of depth of the VP-Tree.
Component ACER(%) AUC(%) EER(%) ACC(%)
16 4.24 95.79 4.26 92.39
20 4.08 96.03 4.03 93.40
29 3.45 96.34 3.44 94.67
40 3.48 97.62 3.90 94.23
60 3.26 98.78 3.67 95.79
TABLE XIII: The impact on the length of prompts.
VI Conclusion
In this paper, we first present a large-scale Unified Face Attack Benchmark (HiFAB), which is the most extensive and comprehensive UAD dataset to date. It includes $14$ kinds of physical attacks, $40$ digital attacks, and $697,347$ videos in total. Every attack is applied to $2,875$ identities to reduce the ID noise. Furthermore, we introduce a novel Hierarchical Prompt Tuning framework (HiPTune), that adaptively integrates classification criteria from various semantic spaces, allowing tailored classification criteria for each sample. Finally, we conduct several experiments that evaluate the model’s performance in real-world scenarios.
VII Disscussion and Future Work
In this work, we explore the application of prompt tuning in the visual domain while acknowledging the value of examining its adaptability to its original text-based purpose. The hierarchical and semantic capabilities of text-based prompt tuning could provide valuable insights for visual applications, warranting further investigation. Furthermore, the suboptimal performance of baseline models under protocol $P_{1}$ may be partially attributed to the absence of a fine-tuning process on the testing datasets. This lack of adaptation limits the models’ ability to generalize to unseen attack types and domain variations. Further discussion on the impact of fine-tuning in bridging this gap will be presented.